OpenStudy (anonymous):

Solve for x: -3(x + 3) = -3(x + 1) - 5.

OpenStudy (rockstar0765):

one question for you what formula do u need or have to use to solve this equation?

OpenStudy (anonymous):

-3(x+3)=-3(x+1)-5 -3x+ -9= -3x+ -3-5 -3x+-9= -3x+-2 +3x+-2=0 x-11=0 x=11 i think this is right and i hope this helps